OpenStudy (anonymous):

Barbara has a part-time lawn-mowing job. The table below shows the amount of money Barbara earned for different numbers of hours of work: Barbara's Earnings from Lawn Mowing Number of Hours Worked Money Earned (dollars) 5 40 6 48 8 56 Part A: What are the missing numbers in the table? Show your work. (4 points)

OpenStudy (anonymous):

Part B: Barbara also works part time at a theme park. The ratio of the number of hours Barbara works at the park to the money she earns in dollars is 1:9. Create a table to show the money earned by Barbara for 5, 6, and 8 hours of work at the theme park. (3 points) Part C: How many more dollars does Barbara earn working at the theme park for 5 hours than mowing lawns for 5 hours? Show your work. (3 points)

OpenStudy (anonymous):

@welshfella @moudown828 I like u too u guys r funny can u help me with another one plzz

OpenStudy (anonymous):

sure

OpenStudy (welshfella):

if you look at the table 40 dollars for 5 hours, 48 dollars for 6 hours how mush is that per hour > 40 / 5 = ? hours

OpenStudy (welshfella):

* dollars not hours ( last word)

OpenStudy (welshfella):

gotta go be back later

OpenStudy (anonymous):

40 / 5 =8

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

@isaac4321 @itsbribro @Phebe @phi @proud_yemeniah_ @Papi_Jamal_5

OpenStudy (anonymous):

can some one help me plzzzzzzzzzzzzzzzzzzzzzzzzzz

OpenStudy (anonymous):

@OnePieceFTW

OpenStudy (onepieceftw):

Alright, so if 40/5 is 8 and 5 * 8 = 40 then the multiplication factor is 8 So what is 8 * 8? Then tell me what is 56/8?

OpenStudy (anonymous):

8*8= 64 56/8= 7

OpenStudy (onepieceftw):

She would be making 8 dollars per hour. part B: the ratio is telling you she is making 9 dollars per hour So just multiply 9 by 5, 6, and 8 Part C: for this just multiply 8 (for the lawn mowing business) by 5 and then 9 (for the amusement park job) by 5

OpenStudy (anonymous):

9*5= 45*6*8= 2160

OpenStudy (anonymous):

@LiveLaughDie

OpenStudy (anonymous):

help me with pt C.